How do you complete the square for the function: y=-0.5x^2+2x+3
Answer by Greenfinch(383)

y=1/2(x^2+ 4x) + 3) make the x^2 coefficient unity
y=1/2(x^2 + 4x +4) -2 +3 create a perfect square and remove the compensation (-2)
y= 1/2(x + 2)^2 +1 y = 0 when x = -2 (+ or -) i